# Tabular foundation model to detect empathy from visual cues

### v1

@@
Abstract.
Detecting empathy from video interactions is an emerging area of research. Video datasets, however, are often released as extracted features (i.e., tabular data) rather than raw footage due to privacy and ethical concerns. Prior research on such tabular datasets established tree-based classical machine learning approaches as the best-performing models. Motivated by the recent success of textual foundation models (i.e., large language models), we explore the use of tabular foundation models in empathy detection from tabular visual features. We experiment with two recent tabular foundation models – TabPFN v2 and TabICL – through in-context learning and fine-tuning setups. Our experiments on a public human-robot interaction benchmark demonstrate a significant boost in cross-subject empathy detection accuracy over several strong baselines (accuracy: $0.590\rightarrow 0.730$; AUC: $0.564\rightarrow 0.669$). In addition to performance improvement, we contribute novel insights and an evaluation setup to ensure generalisation on unseen subjects in this public benchmark. As the practice of releasing video features as tabular datasets is likely to persist due to privacy constraints, our findings will be widely applicable to future empathy detection video datasets as well.
@@ Abstract.
empathy, tabular data, foundation model, empathy detection, human-robot interaction, video features
ccs: Computing methodologies Learning paradigmsccs: Applied computing Psychologyccs: Computing methodologies Neural networks
1. Introduction
@@
Figure 1. Overview of the task and motivation. We study the task of detecting empathy of subjects listening to narrative storytelling, where the listener self-reports their empathy level. From the listener’s video, visual features (e.g., eye gaze, facial expressions) are released as a tabular dataset (Mathur et al., 2021). In this context, we propose using tabular foundation models. Our results show that both in-context learning and fine-tuning approaches improve performance, and fine-tuning provides stronger generalisation to unseen subjects.
@@
Empathy, as defined in social psychology (Davis, 1996; Hoffman, 2000), refers to an individual’s capacity to perceive, comprehend and vicariously experience the emotional state of others. This interpersonal skill has practical significance in a variety of interaction contexts, such as human-robot interactions, clinical communication and educational settings. Training intervention to nurture this skill requires measurement [^1] We use the terms ‘measurement’, ‘detection’ and ‘prediction interchangeably. of empathy. Current assessment approaches rely on self-reported questionnaires or labour-intensive manual analysis by external observers (Bas-Sarmiento et al., 2020). Automated empathy detection through machine learning techniques has gained momentum over the past decade (Hasan et al., 2024b). While the majority of this progress has occurred within natural language processing contexts, i.e., detecting empathy in textual content, empathy detection from video data remains relatively under-explored (Hasan et al., 2024b). Towards this direction, this paper measures empathy from video-based behavioural cues.
@@
Several empathy detection benchmarks are proposed in the literature aiming at different aspects of empathy, such as assessing the level of empathy in an interaction (Buechel et al., 2018; Mathur et al., 2021), measuring empathic similarity between two contents (Shen et al., 2023), measuring the changes of someone’s emotional state (Barros et al., 2019), and detecting the direction of empathy (Hosseini and Caragea, 2021). Our work is grounded on the assessment of the level of empathy, being the most common form in the literature (Hasan et al., 2024b). In particular, we detect the level of empathy elicited by storytelling, i.e., measuring how much empathy a listener demonstrates towards others portrayed in a story (Mathur et al., 2021; Keen, 2006) (Figure 1).
@@
Being an emerging area of research, the amount of public video datasets is very limited (Hasan et al., 2024b). Datasets of raw video footage featuring people’s activity are difficult to make public due to privacy and ethical concerns. Consequently, the majority of these few existing public datasets, such as human-robot interactions (Mathur et al., 2021; Spitale et al., 2022), are released as extracted feature vectors (i.e., tabular data) instead of raw video footage. Since privacy and ethical concerns are likely to persist, one of the key challenges in developing an empathy detection system from video is machine learning from tabular datasets.
@@
Traditional approaches to tabular data have long been dominated by tree-based methods like gradient-boosted decision trees (Hollmann et al., 2023, 2025). While deep learning has revolutionised learning from texts and images, it has struggled with tabular data due to the heterogeneity between datasets and the heterogeneous nature of the raw data itself. Foundation models like large language models (LLMs) and large vision models (LVMs) have been widely explored in downstream tasks on texts and videos. Similarly, LLMs have also been explored with tabular data (Hegselmann et al., 2023; Fang et al., 2024), but these approaches convert table rows into some textual format, such as sentences. As LLMs do not natively support numerical tabular data, a new paradigm of the foundation model has emerged: the tabular foundation model. The first one of its kind, a prior data-fitted network (PFN) (Müller et al., 2022) trained on lots of synthetic tabular data, is TabPFN (Hollmann et al., 2023). Following the success of TabPFN, TabICL (tabular in-context learning) (Qu et al., 2025) is another tabular foundation model, introduced in 2025.
@@
Among the few video-based public empathy detection benchmarks, Mathur et al. (2021) features empathy elicited by a robot storyteller. In this dataset, human subjects attend to a robot telling different stories that may elicit empathy in the human subjects. Although there are multiple sessions per subject, prior work has failed to consider scenarios where the same subject’s data may appear in both training and test sets. Given that empathy is subjective, the same subjects appearing in both training and test splits do not ensure true generalisation performance on unseen subjects. To address this, we design experiments and report results in a new evaluation scheme on this benchmark dataset where it is ensured that the same subject’s data is not shared between the training and test sets. Our proposed fine-tuning of the tabular foundation model achieves impressive results on this cross-subject generalisation.
Previous work on video-based elicited empathy detection demonstrated classical machine learning models as state-of-the-art approaches. In contrast, we explore the potential of tabular foundation models and outperform existing methods by a significant margin (Section 4.3 and Section 4.4). Our key contributions are threefold:
  1. (1)
    We design and evaluate in-context learning (ICL) with two tabular foundation models (TabPFN and TabICL) for empathy detection from visual cues.
  2. (2)
@@
    Our proposed fine-tuning of the popular TabPFN foundation model results in further performance improvement, especially in detecting empathy of unseen subjects.
  3. (3)
@@
    We reveal limitations in current evaluation protocols in a public empathy detection benchmark and propose a fairer and unbiased evaluation to ensure robust empathy detection.
2. Related work
2.1. Empathy detection
@@
There are several empathy detection studies in text domains, such as written essays (Buechel et al., 2018; Hasan et al., 2023), social media conversations (Khanpour et al., 2017) and email correspondence (Sedefoglu et al., 2024). In detecting empathy from videos, Mathur et al. (2021) released a human-robot interaction (HRI) dataset, aiming to measure empathy elicited by a robot storyteller. They evaluated 10 different machine learning approaches, including eight classical models (e.g., AdaBoost, bagging, decision trees, support vector machine, logistic regression, random forest and XGBoost) and two deep learning architectures (long short-term memory network and temporal convolutional network). For classical machine learning approaches, features were represented as fixed-length vectors of statistical attributes (mean, median, standard deviation and autocorrelation). For deep learning approaches, the features were maintained as temporal sequences. Through 5-fold stratified cross-validation with 10 repeats, XGBoost appeared as the highest-performing classifier. Contrary to their work, we propose in-context learning and fine-tuning of tabular foundation models in detecting elicited empathy.
@@
Other studies on video-based empathy detection include interaction between human subjects and virtual agents (Kroes et al., 2022; Tavabi et al., 2019), teacher and students (Pan et al., 2022), human subject and avatar (Hervás et al., 2016), counsellor and clients (Zhu et al., 2023), two human subjects (Barros et al., 2019), and human and AI agents (Shen et al., 2024). In contrast to these empathy modelling approaches, we focus on measuring elicited empathy from the public HRI benchmark.
2.2. Application of foundation models
@@
Textual foundation models (i.e., LLMs) have been explored in several stages of empathy detection from textual data. For example, Hasan et al. (2025) used GPT-4 and Llama-3 LLMs and Kong and Moon (2024) used GPT series LLMs in a zero-shot setting to predict empathy in written essays. On the same task, Kong and Moon (2024) and Hasan et al. (2024a) explored few-shot prompt engineering (i.e., in-context learning) with GPT LLMs. Apart from the prediction, LLMs are also explored in data augmentation (Lu et al., 2023; Frick and Steinebach, 2024) and data annotation (Hasan et al., 2025, 2024a).
@@
Motivated by the success of LLMs in empathy detection from text, we explore two recent tabular foundation models released in 2025 – TabPFN v2 (Hollmann et al., 2025) and TabICL (Qu et al., 2025) – for tabular video features. While tabular foundation models’ in-context learning has been demonstrated in numerical datasets such as OpenML test suites (Vanschoren et al., 2014), we explore its capability in dealing with video features. In addition to the in-context learning, we fine-tune TabPFN v2, which further boosts empathy classification performance.
3. Method
@@
Figure 2. Overview of the proposed pipeline for empathy detection using tabular foundation models. Temporal multimodal features (e.g., facial landmarks, eye gaze) are first extracted for each sample $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$. These sequences are summarised by computing $n$ statistical attributes (e.g., mean, median) over time, resulting in a fixed-size tabular representation $\mathbf{X}_{i}\in\mathbb{R}^{n\times d}$. This representation is used for binary empathy classification $y_{i}\in\{0,1\}$) via either (top) in-context learning, where a frozen tabular foundation model predicts labels with a single forward pass over the training data, or (bottom) fine-tuning of the foundation model with frozen X and Y encoders. The model predicts whether a subject is empathic or less empathic.
@@
Figure 2 presents an overview of our approach for empathy detection using tabular foundation models. Temporal multimodal features, such as facial landmarks and eye gaze, are first aggregated into fixed-size tabular representations via statistical summarisation. These are then used for binary empathy classification through either in-context learning (ICL) or fine-tuning of a tabular foundation model. Further details of each component are provided in the following subsections.
3.1. Problem formulation
@@
We formulate empathy detection in human-robot interaction videos as a supervised or in-context learning problem. Given a dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$, consisting of $N$ interaction samples from $M$ subjects, where $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$ represents a $d$-dimensional feature vector across $T_{i}$ frames of the $i$-th video sample and $y_{i}\in\{0,1\}$ denotes the corresponding binary empathy label.
@@
The feature vector $\mathbf{X}_{i}$ contains frame-level visual behavioural cues (e.g., eye gaze directions, facial action unit intensities, facial landmarks, and other non-verbal indicators) extracted using toolkits such as OpenFace (Amos et al., 2016). These temporal features can be processed in two ways: (1) time-series modelling or (2) transformed into fixed-length representations by computing statistical attributes (e.g., mean, median, standard deviation and autocorrelation) across the temporal axis. Prior work (Mathur et al., 2021) demonstrated that the latter approach works better for empathy detection in the HRI context.
@@
For $n$ statistical attributes, the input feature representation becomes a fixed-size representation $\mathcal{X}_{i}\in\mathbb{R}^{n\times d}$. To reduce dimensionality, we select $d^{\prime}$ features from the original $d$ features using ANOVA F-values computed between each feature and the target labels. Given this tabular representation, our goal is to learn a function $f:\mathbb{R}^{n\times d^{\prime}}\rightarrow[0,1]$ that maps the input features to a conditional probability distribution over the target variable (empathy level). For a general supervised learning task, it becomes:
(1) $p(y=1|\mathbf{x};\theta)=f_{\theta}(\mathbf{x})$
@@
3.2. Traditional supervised learning approaches
@@
(2) $\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta,\mathcal{D}_{\text{train}})$
where $\mathcal{L}$ is a loss function (e.g., binary cross-entropy) and $\theta^{*}$ represents the optimised parameters. During inference, predictions are made on a separate test data:
(3) $p(y=1|\mathbf{x};\theta^{*})=f_{\theta^{*}}(\mathbf{x})$
3.3. In-context learning with tabular foundation model

### v2

# TFMPathy: Tabular Foundation Model for Privacy-Aware, Generalisable Empathy Detection from Videos

@@
Abstract
Detecting empathy from video interactions is an emerging area of research, particularly in healthcare and social robotics. However, privacy and ethical concerns often prevent the release of raw video data, with many datasets instead shared as pre-extracted tabular features. Previous work on such datasets has established classical tree-based models as the state of the art. Motivated by recent successes of large-scale foundation models for text, we investigate the potential of tabular foundation models (TFMs) for empathy detection from video-derived tabular data. Our proposed system, TFMPathy, is demonstrated with two recent TFMs (TabPFN v2 and TabICL) under both in-context learning and fine-tuning paradigms. On a public human-robot interaction benchmark, TFMPathy significantly improves empathy detection accuracy reported in the literature. While the established evaluation protocol in the literature does not ensure cross-subject generalisation, our evaluation scheme also captures such generalisation. We show that TFMPathy under a fine-tuning setup has better cross-subject generalisation capacity over baseline methods (accuracy: $0.590\rightarrow 0.730$; AUC: $0.564\rightarrow 0.669$). Given the ongoing privacy and ethical constraints around raw video sharing, the proposed TFMPathy system provides a practical and scalable path toward building AI systems dependent on human-centred video datasets. Our code is publicly available at https://github.com/hasan-rakibul/TFMPathy (will be made available upon acceptance of this paper).
Index Terms:
@@ Abstract
Empathy detection, human-robot interaction, tabular data, foundation model, video features, privacy, generalisation.
publicationid: pubid: This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice.
I Introduction
@@
Figure 1: Overview of the task and motivation. We study the task of detecting empathy of subjects listening to narrative storytelling, where the listener self-reports their empathy level. From the listener’s video, visual features (e.g., eye gaze, facial expressions) are released as a tabular dataset [1]. In this context, we propose TFMPathy, which uses tabular foundation models (TFMs) as the backbone. Our results show that TFMPathy in both in-context learning (ICL) and fine-tuning approaches improves performance, and fine-tuning provides stronger generalisation to unseen subjects.
@@
Empathy, as defined in social psychology [2, 3], refers to an individual’s capacity to perceive, comprehend and vicariously experience the emotional state of others. This interpersonal skill has practical significance in a variety of interaction contexts, such as human-robot interactions (HRIs), clinical communication and educational settings. Training intervention to nurture this skill requires measurement [^1] We use the terms ‘measurement’, ‘detection’ and ‘prediction’ interchangeably. of empathy. Current assessment approaches rely on self-reported questionnaires or labour-intensive manual analysis by external observers [4]. Automated empathy detection through machine learning techniques has gained momentum over the past decade [5]. While the majority of this progress has occurred within natural language processing contexts, i.e., detecting empathy in textual content, empathy detection from video data remains relatively under-explored [5]. Towards this direction, this paper measures empathy from video-based behavioural cues.
@@
Several empathy detection benchmarks are proposed in the literature aiming at different aspects of empathy, such as assessing the level of empathy in an interaction [6, 1], measuring empathic similarity between two contents [7], measuring the changes of someone’s emotional state [8], and detecting the direction of empathy [9]. Our work is grounded on the assessment of the level of empathy, being the most common form in the literature [5]. In particular, we detect the level of empathy elicited by storytelling, i.e., measuring how much empathy a listener demonstrates towards others portrayed in a story [1, 10] (Figure 1).
@@
Being an emerging area of research, the amount of public video datasets is very limited [5]. Datasets of raw video footage featuring people’s activity are difficult to make public due to privacy and ethical concerns. Consequently, the majority of these few existing public datasets, such as human-robot interactions [1, 11], are released as extracted feature vectors (i.e., tabular data) instead of raw video footage. Since privacy and ethical concerns are likely to persist, one of the key challenges in developing an empathy detection system from video is machine learning from tabular datasets.
@@
Traditional approaches to tabular data have long been dominated by tree-based methods like gradient-boosted decision trees [12, 13]. While deep learning has revolutionised learning from texts and images, it has struggled with tabular data due to the heterogeneity between datasets and the heterogeneous nature of the raw data itself. Foundation models like large language models (LLMs) and large vision models (LVMs) have been widely explored in downstream tasks on texts and videos. Similarly, LLMs have also been explored with tabular data [14, 15], but these approaches convert table rows into textual formats such as sentences. As LLMs do not natively support numerical tabular data, a new paradigm of the foundation model has emerged: the tabular foundation model (TFM) [^2] Not to be confused with LLMs. While both are pre-trained foundation models, TFMs are trained with numerical tabular data, whereas LLMs are trained with texts. TFMs are significantly smaller than LLMs. For example, TabPFN has only 7.2 million trainable parameters, in contrast to billions in typical LLMs.. The first one of its kind, a prior data-fitted network (PFN) [16] trained on lots of synthetic tabular data, is TabPFN [12]. Following the success of TabPFN, TabICL (tabular in-context learning) [17] is another TFM, introduced recently in 2025. TFMs natively support in-context learning (ICL), which is a paradigm in which a pre-trained foundation model is conditioned on a small set of input-output examples at inference time. With these examples, the model adapts to new tasks without any gradient updates (i.e., training).
@@
Among the few video-based public empathy detection benchmarks, [1] features empathy elicited by a robot storyteller. In this HRI dataset, human subjects attend to a robot telling different stories that may elicit empathy in the human subjects. Although there are multiple sessions per subject, prior work has failed to consider scenarios where the same subject’s data may appear in both training and test sets. Given that empathy is subjective, such overlaps do not ensure true generalisation performance on unseen subjects. To address this, we design experiments and report results in a new evaluation scheme on this benchmark dataset, where it is ensured that the same subject’s data is not shared between the training and test sets. Our proposed fine-tuning of the TFM achieves impressive results on this cross-subject generalisation (Section IV-D).
Novelty and contributions
Previous work on video-based elicited empathy detection demonstrated classical machine learning models as state-of-the-art approaches. In contrast, our proposed TFMPathy system explores the potential of TFMs and outperforms existing methods by a significant margin (Section IV-C and Section IV-D). We are the first to fine-tune TabPFN to video-based empathy detection with privacy constraints. To make fine-tuning effective and practical, we address key challenges such as catastrophic forgetting and conduct extensive ablation studies covering layer freezing, optimiser and scheduler choices, batch size, softmax temperature and feature dimensionality (Section IV-E). While TFMs are designed for numerical tabular data, we extend them to visual features, in a similar spirit to many works on using LLMs on new domains [18, 19]. We provide new insights and practical strategies for adapting TFMs to privacy-constrained, real-world data derived from videos. On the public HRI dataset, we establish a cross-subject generalisation benchmark, which has not been attempted before. Our key contributions are threefold:
  1. 1.
    We explore the novel application of two TFMs (TabPFN and TabICL) in the ICL setup for empathy detection from video-derived visual features under privacy constraints.
  2. 2.
@@
    Our proposed TFMPathy system introduces a practical and effective fine-tuning strategy for the popular TabPFN foundation model for empathy detection, which demonstrates further performance improvement than ICL, especially in detecting the empathy of unseen subjects.
  3. 3.
@@
    We enhance current evaluation protocols of a public empathy detection benchmark by introducing a subject-independent evaluation benchmark to ensure cross-subject generalisation and to enable fairer and unbiased assessment for real-world applicability.
II Related work
II-A Empathy detection
@@
There are several empathy detection studies in text domains, such as written essays [6, 20], social media conversations [21] and email correspondence [22]. In detecting empathy from videos, [1] released a human-robot interaction (HRI) dataset, aiming to measure empathy elicited by a robot storyteller. They evaluated 10 different machine learning approaches, including eight classical models (e.g., AdaBoost, bagging, decision trees, support vector machine, logistic regression, random forest and XGBoost) and two deep learning architectures (long short-term memory network and temporal convolutional network). For classical machine learning approaches, features were represented as fixed-length vectors of statistical attributes (mean, median, standard deviation and autocorrelation). For deep learning approaches, the features were maintained as temporal sequences. Through 5-fold stratified cross-validation with 10 repeats, XGBoost appeared as the highest-performing classifier. Contrary to their work, we propose ICL and fine-tuning of TFMs in detecting elicited empathy.
@@
Other studies on video-based empathy detection include interaction between human subjects and virtual agents [23, 24], teacher and students [25], human subject and avatar [26], counsellor and clients [27], two human subjects [8], and human and AI agents [28]. In contrast to these empathy modelling approaches, we focus on measuring elicited empathy from the public HRI benchmark.
II-B Application of foundation models
@@
Textual foundation models (i.e., LLMs) have been explored in several stages of empathy detection from textual data. For example, [29] used GPT-4 and Llama-3 LLMs, and [30] used GPT series LLMs in a zero-shot setting to predict empathy in written essays. On the same task, [30] and [31] explored few-shot prompt engineering (i.e., ICL) with GPT LLMs. Apart from the prediction, LLMs are also explored in data augmentation [32, 33] and data annotation [29, 31].
@@
Motivated by the success of LLMs in empathy detection from text, we explore two recent TFMs released in 2025 – TabPFN v2 [13] and TabICL [17] – for tabular video features. While TFMs’ ICL capacity has been demonstrated in numerical datasets such as OpenML test suites [34], we explore its capability in dealing with video features. In addition to ICL, we fine-tune TabPFN v2, which further boosts empathy classification performance.
III Method
@@
Figure 2: Overview of our proposed TFMPathy pipeline for empathy detection. Temporal multimodal features (e.g., facial landmarks, eye gaze) are first extracted for each sample $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$. These sequences are summarised by computing $n$ statistical attributes (e.g., mean, median) over time, resulting in a fixed-size tabular representation $\mathbf{X}_{i}\in\mathbb{R}^{n\times d}$. This representation is used for binary empathy classification $y_{i}\in\{0,1\}$) via either (top) in-context learning, where a frozen tabular foundation model (TFM) predicts labels with a single forward pass over the training data, or (bottom) fine-tuning of the foundation model with frozen X and Y encoders. The model predicts whether a subject is empathic or less empathic.
@@
Figure 2 presents an overview of our approach for empathy detection using TFMs. Temporal multimodal features, such as facial landmarks and eye gaze, are first aggregated into fixed-size tabular representations via summary statistics. These are then used for binary empathy classification through either ICL or fine-tuning of a TFM. Further details of each component are provided in the following subsections.
III-A Problem formulation
@@
We formulate empathy detection in HRI videos either as a supervised learning or an ICL problem. Given a dataset $\mathcal{D}=\{(\mathbf{X}_{i},y_{i})\}_{i=1}^{N}$, consisting of $N$ interaction samples from $M$ subjects, where $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$ represents a $d$-dimensional feature vector across $T_{i}$ frames of the $i$-th video sample and $y_{i}\in\{0,1\}$ denotes the corresponding binary empathy label.
III-B Data preprocessing for tabular foundation model
@@
The feature vector $\mathbf{X}_{i}$ contains frame-level visual behavioural cues (e.g., eye gaze directions, facial action unit intensities, facial landmarks, and other non-verbal indicators) extracted using toolkits such as OpenFace [35]. These temporal features can be processed in two ways: (1) time-series modelling or (2) transformed into fixed-length representations by computing statistical attributes (e.g., mean, median, standard deviation and autocorrelation) across the temporal axis. Prior work [1] demonstrated that the latter approach works better for empathy detection in the HRI context.
From $\mathbf{X}_{i}\in\mathbb{R}^{T_{i}\times d}$, we first filter out temporal points having zeros in all features, and then calculate statistical attributes. For $n$ statistical attributes, the input feature representation becomes a fixed-size representation $\mathcal{X}_{i}\in\mathbb{R}^{nd}$. For reference, in OpenFace features (709 dimensions) and four statistical attributes, the resultant feature becomes a vector of $709\times 4=2836$ elements. From this, we first drop dimensions where it is constant across all samples.
Both TabPFN and TabICL support up to 500 features, and therefore, we further reduce the feature dimension. We rank all features by their ANOVA F‑values with the target labels on the training set and retain the top $d^{\prime}$ features. At evaluation time, we apply this fixed subset to each test split, without recomputing F‑values to prevent any information leakage from test to training. This feature selection process is performed independently within each fold of cross-validation.
@@
With this tabular representation, our goal is to learn a function $f:\mathbb{R}^{d^{\prime}}\rightarrow[0,1]$ that maps the input features to a conditional probability distribution over the target variable (empathy level). For a general supervised learning task, it becomes:
$p(y=1|\mathbf{x};\theta)=f_{\theta}(\mathbf{x})$ (1)
@@
III-C Traditional supervised learning approaches
@@
$\theta^{*}=\arg\min_{\theta}\mathcal{L}(\theta,\mathcal{D}_{\text{train}})$ (2)
where $\mathcal{L}$ is a loss function (e.g., binary cross-entropy) and $\theta^{*}$ represents the optimised parameters. During inference, predictions are made on a separate test set:
$p(y=1|\mathbf{x};\theta^{*})=f_{\theta^{*}}(\mathbf{x})$ (3)
III-D In-context learning with tabular foundation model